OpenStudy (anonymous):

Verify the identity. sec x + tan x = cos x/ 1-sin x

OpenStudy (anonymous):

sec = 1/cos and tan = sin/cos so lhs is (1 +sin)/cos Cross multiply 1 - sin^2 = cos^2 and Cos^2 +Sin^2 = 1 QED